OpenStudy (anonymous):

solve the inequality. graph the solution set. 5r+4<=5 r<= 1/5 r<= 1 4/5 r>= 1/5 r>= 1 4/5

OpenStudy (anonymous):

5R - 1/5 = 4/5 5R = 4/5 + 1/5 R = 1/5

OpenStudy (anonymous):

I think it's C

OpenStudy (anonymous):

so its either a or c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

It definatley C

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

did you get the answer